OpenStudy (anonymous):

what is 0/0 equal to ?

OpenStudy (ksaimouli):

o orf undefined

OpenStudy (anonymous):

undefined if you are solving a limit you have made a mistake and can simplify

OpenStudy (anonymous):

end of the world. (run)

OpenStudy (anonymous):

haha joe

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

it depends on what you are asking here If you are saying to the 0th power: (3ab)^0 = 1 If you are multiplying by 0: (3ab) * 0 = 0

OpenStudy (anonymous):

he is talking about division. The question was what is: \[\frac{0}{0}\]and the answer is undefined. The reason its undefined is because we want to preserve the axiom that if "a" is any number, then: \[a\cdot 0 = 0\] If we allowed division by 0, then there would be some number, say alpha, such that: \[\alpha =\frac{a}{0} \]What would we get when we multiply alpha by 0? Just like: \[\frac{1}{3}\cdot 3=1, \frac{4}{5}\cdot5=4, \frac{123}{45}\cdot45=123\] we would get: \[\alpha \cdot0=\frac{a}{0}\cdot 0=a\] So alpha times 0 equals a, which contradicts the axiom that any number times 0 equals 0. This is why we cant divide by 0.